# SketchSplat: 3D Edge Reconstruction via Differentiable Multi-view Sketch Splatting

### v1

@@ Abstract
Edges are one of the most basic parametric primitives to describe structural information in 3D. In this paper, we study parametric 3D edge reconstruction from calibrated multi-view images. Previous methods usually reconstruct a 3D edge point set from multi-view 2D edge images, and then fit 3D edges to the point set. However, noise in the point set may cause gaps among fitted edges, and the recovered edges may not align with input multi-view images since the edge fitting depends only on the reconstructed 3D point set. To mitigate these problems, we propose SketchSplat, a method to reconstruct accurate, complete, and compact 3D edges via differentiable multi-view sketch splatting. We represent 3D edges as sketches, which are parametric lines and curves defined by attributes including control points, scales, and opacity. During edge reconstruction, we iteratively sample Gaussian points from a set of sketches and rasterize the Gaussians onto 2D edge images. Then the gradient of the image error with respect to the input 2D edge images can be back-propagated to optimize the sketch attributes. Our method bridges 2D edge images and 3D edges in a differentiable manner, which ensures that 3D edges align well with 2D images and leads to accurate and complete results. We also propose a series of adaptive topological operations and apply them along with the sketch optimization. The topological operations help reduce the number of sketches required while ensuring high accuracy, yielding a more compact reconstruction. Finally, we contribute an accurate 2D edge detector that improves the performance of both ours and existing methods. Experiments show that our method achieves state-of-the-art accuracy, completeness, and compactness on a benchmark CAD dataset.
@@ 1 Introduction
Traditional methods [12, 5, 47, 23] usually detect 2D line segments in each view. Then the 2D lines are lifted to 3D via line matching and triangulation across views. While these methods achieve high accuracy and efficiency, robust line matching and consistent detection across views are difficult to achieve, leading to incomplete results. In addition, these methods are usually not capable of handling lines and curves simultaneously.
In recent years, differentiable rendering has emerged as a promising technique to reconstruct 3D content for novel view synthesis [27, 3, 28, 54], surface reconstruction [46, 25, 22, 11], and animation [29, 34, 48]. In the field of edge reconstruction, several works introduced the idea of differentiable rendering to recover edge information [53, 19, 7]. Instead of detecting parametric 2D edge segments explicitly, they leverage pixel-wise 2D edge intensity images [43, 33] as a supervision signal and optimize a point-based 3D edge representation via differentiable multi-view rendering, resulting in an extracted 3D edge point set. Parametric 3D edges are then obtained by fitting lines and curves to the extracted 3D points. Since the edge-fitting step relies purely on the recovered 3D edge points but not the input 2D edge images, the fitted 3D edges may not align well with the 2D edge images. In addition, noise in the 3D edge points usually leads to discontinuous and fragmented edges.
To address these limitations, we propose SketchSplatting, a differentiable edge reconstruction framework that directly optimizes parametric 3D edges by fitting them to the input 2D edge images using differentiable rendering. Our experiments show that this leads to more accurate edge reconstruction with a compact set of parametric edges that more faithfully capture the original connectivity.
@@ 2 Related Work
The advancement of differentiable rendering opens up a new way for 3D reconstruction [27, 46, 22]. Inspired by NeRF [27], NEF [53] proposes to optimize an implicit neural radiance field with a multi-view rendering loss. After training, 3D edge points are extracted from the radiance field and curves are fitted to these points. NEAT [50] builds on VolSDF [51] and uses a signed distance field to represent both object surfaces and endpoints of 3D lines. The surface-based modeling approach restricts this method to objects with rich texture. EMAP [19] adopts an unsigned distance function (UDF), where sampled points can be displaced towards the zero-level set of the UDF to be more accurate. While methods based on neural field are time consuming to optimize, EdgeGS [7] represents 3D edge fields with explicit Gaussian points and leverages Gaussian splatting [17] to render and optimize Gaussian points, leading to more efficient edge reconstruction. A common issue of these methods is that the edge fitting process relies only on the extracted 3D edge points, hence noise in the reconstructed 3D points may lead to inaccurate and incomplete edges. Instead we represent 3D edges as sketches that are defined parametrically using control points, scales and opacities, and directly optimize the sketch parameters using differentiable rendering. Differentiable rendering uses sketch sampling and rasterization to align the 3D edges with 2D edge images directly, yielding more accurate and complete results.
@@
DiffVG [21] proposes a differentiable rasterizer that can optimize 2D vector graphics with 2D image supervision in a differentiable manner. However, the differentiation process relies on distance measurements between the rasterized curves and pixels, which needs to be custom tailored for each type of curve. 3Doodle [9] extends the idea of DiffVG [21] to reconstruct 3D curve-based object abstractions using multiple input views. They optimize a set of Bézier curves by minimizing the perceptual loss between the rasterized image and multi-view CLIP feature maps [37, 44]. Since CLIP features only contain abstract information, 3Doodle is not suitable for recovering accurate edges of objects like CAD models. They also approximate the projection of Bézier curves using an orthographic camera model. Instead, we get inspiration from 3DGS [17] and propose to optimize 3D sketches via sampling and rasterization of Gaussian points sampled from sketch parameters, which can be easily extended to different types of curves and achieves accurate 3D edge reconstruction.
@@ 4 Method
Given multi-view posed RGB images as input, our target is to recover the parametric edges for the 3D objects. Fig. 2 provides an overview of our approach. We firstly compute an edge image for each RGB image. Then we represent edges as 3D parametric sketches, including both lines $l\in\mathbb{R}^{2\times 3}$ and third-order Bézier curves with four 3D control points $c\in\mathbb{R}^{4\times 3}$. We initialize a set of sketches from existing edge reconstruction methods, and we set the control points of sketches as optimizable parameters. During sketch optimization, we use a sample-based method to convert sketches to Gaussian points, which can be rasterized onto 2D images to calculate the image loss. We back-propagate the gradient of the loss to the sketch parameters to optimize them using gradient descent. In addition, we introduce a set of topological operations to adaptively refine the sketches during training, yielding more accurate and compact edge reconstruction results.
Figure 2: Overview of SketchSplat pipeline: (a) We extract edge images using a novel approach (Section 4.1). (b) We then obtain initial edge points using EdgeGS [7] and (c) fit parametric sketches defined as lines and third-order Bézier curves. The core step of our approach is to optimize the sketches by (d) sampling and rasterizing them in a differentiable manner and back-propagating gradients of the image loss to (e) sketch parameters. In addition, during training we apply (f) a set of topological operations to improve compactness and connectivity of the reconstructed sketches. (g) After training, we take the optimized geometric parameters of sketches as the reconstructed 3D edges.
@@ 4.1 Edge Image Extraction
Previously, some methods [23, 50] extract edges with line detectors like LSD [45] and SOLD2 [30]. While these methods produce accurate parametric lines in 2D, they may fail to handle curve regions and result in incomplete results. Recent methods [53, 19, 7] use neural network-based edge detectors [43, 33] to detect edge intensity maps as supervision signal. However, since these detectors are not trained specifically on CAD-like objects, they may miss some significant edges in 2D or wrongly recover shading textures as real edges. Finally, we observed that these detectors introduce bias by consistently shifting detected edges slightly from certain real edges, which leads to multi-view inconsistency in 3D edge reconstruction. We address these issues by proposing a new edge detection method to mitigate the mentioned problems. In particular, our approach leverages the fact that our problem statement assumes the availability of multiple input images.
For each RGB image, we combine information from three parts: foreground mask $A$, depth map $D$, and normal map $N$. The foreground mask can be extracted using the image alpha-channel, a thresholding operation or possibly using semantic segmentation, providing object boundary information. We obtain a depth image for each view by reconstructing the scene with 2DGS [14] and rendering depth maps from each view. We use a normal estimator [52] to generate normal maps $N$ for each view. Then we calculate gradient maps $g(D)$ and $g(N)$ with the Sobel operator [42], threshold them with $t_{d}$ and $t_{n}$, and combine them via the OR operation, followed by a Gaussian filter $G_{f}(\cdot)$ to smooth the boundaries of the edges. In summary, we call our edge detector 2DGS-SN, defined as
@@ 4.2 Sketch Splatting
We represent edges as 3D parametric sketches, including both lines $l\in\mathbb{R}^{2\times 3}$ and third-order Bézier curves with control points $c\in\mathbb{R}^{4\times 3}$. This is the same as the output format of EdgeGS[7]. While EdgeGS optimizes 3D point positions first and then fits 3D lines and curves to the 3D points, we optimize the parameters of lines and curves directly via multi-view differentiable rasterization. In addition to the optimizable geometry parameters ($l$ or $c$), we equip each sketch with an optimizable opacity $o\in\mathbb{R}$ and local scale $s\in\mathbb{R}^{3}$ to model the sketch intensity and thickness, respectively. These attributes will support Gaussian sampling and rasterization in the next steps.
@@
Our key idea is to bridge the gap between 2D edge images and 3D sketches via gradient descent optimization and differentiable rendering, which significantly improves sketch reconstruction over the initialization using EdgeGS [7]. Inspired by 3DGS [17] and DiffVG [21], we choose 3D Gaussians as an intermediate representation to achieve differentiable sketch rendering.
For each gradient descent step, we first sample points from each sketch with a constant step size (5mm for all experiments), resulting in a point set $P$. Each point will be assigned the opacity and scale value of the corresponding sketch. To model the orientation, the main direction of each point is sampled as the local tangent direction of the sketch. Then we follow 3DGS[17] to rasterize the Gaussian points onto a chosen 2D view to generate a rendered edge image $E^{*}$. We calculate the $L_{1}$ loss between $E^{*}$ and the ground truth edge image $E_{gt}$, and back-propagate the gradient to update the sketch parameters $S_{i}=\{l_{i}\ \text{or}\ c_{i},o_{i},s_{i}\}$.
@@ 4.3 Adaptive Topology Control
In our representation, each sketch has two end-points. If there exist two end-points, one from sketch $S_{i}$ and one from sketch $S_{j}$, that are sufficiently close (i.e. the point distance is smaller than threshold $th_{connect}=10mm$), we will connect sketch $S_{i}$ and $S_{j}$ by merging those two end-points into one point (see Fig. 2(f-i)). However, this means the number of optimizable points differs from the number of control points |$|\{l\}|*2+|\{c\}|*4$. To accommodate this, we maintain an optimizable point set $P$, and the control points of each line $l$ and curve $c$ are represented by indexing point set $P$. This representation also supports our additional topological operations.
@@
As shown in Fig. 2(f-ii), For two sketches, if one of them can be almost completely covered by the other one, the smaller one should be merged into the larger one. Specifically, for sketch $S_{i}$ and $S_{j}$, we sample two point sets $P_{i}$ and $P_{j}$ and calculate the distance between each pair of points. Then for $P_{i}$, we calculate the ratio $r_{i\rightarrow j}$ of points that have close neighboring points in $P_{j}$, where the neighbor threshold is $th_{neighbor}=10mm$. If $r_{i\rightarrow j}>th_{overlap}=80\%$, sketch $S_{i}$ will be merged in sketch $S_{j}$. If both $r_{i\rightarrow j}$ and $r_{j\rightarrow i}$ surpass $th_{overlap}$, we will choose the sketch with larger overlapping ratio to be merged.
@@ 4.4 Implementation Details
Unless otherwise specified, we use EdgeGS [7] as our default initialization method, as it provides a sufficient number of 3D edges to represent the complete structure of 3D objects. For each scene, we optimize the sketches for 1000 epochs with Adam. We update parameters only once for each epoch, where the loss is accumulated across all the training views. All our experiments are implemented on a single RTX A5000 NVIDIA GPU. In average, each scene costs about 10 minutes for training, where both initialization and sketch optimization takes about 5 minutes, respectively.

### v2

@@ Abstract
Edges are one of the most basic parametric primitives to describe structural information in 3D. In this paper, we study parametric 3D edge reconstruction from calibrated multi-view images. Previous methods usually reconstruct a 3D edge point set from multi-view 2D edge images, and then fit 3D edges to the point set. However, noise in the point set may cause gaps among fitted edges, and the recovered edges may not align with input multi-view images since the edge fitting depends only on the reconstructed 3D point set. To mitigate these problems, we propose SketchSplat, a method to reconstruct accurate, complete, and compact 3D edges via differentiable multi-view sketch splatting. We represent 3D edges as sketches, which are parametric lines and curves defined by attributes including control points, scales, and opacity. During reconstruction, we iteratively sample Gaussian points from a set of sketches and rasterize the Gaussians onto 2D edge images. Then the gradient of the image loss can be back-propagated to optimize the sketch attributes. Our method bridges 2D edge images and 3D edges in a differentiable manner, which ensures that 3D edges align well with 2D images and leads to accurate and complete results. We also propose a series of adaptive topological operations to reduce redundant edges and apply them along with the sketch optimization, yielding a more compact reconstruction. Finally, we contribute an accurate 2D edge detector that improves the performance of both ours and existing methods. Experiments show that our method achieves state-of-the-art accuracy, completeness, and compactness on a benchmark CAD dataset.
@@ 1 Introduction
Traditional methods [13, 5, 49, 24] usually detect 2D line segments in each view. Then the 2D lines are lifted to 3D via line matching and triangulation across views. While these methods achieve high accuracy and efficiency, multi-view consistent detection and robust line matching are difficult to achieve, leading to incomplete results. In addition, these methods are usually not capable of handling lines and curves simultaneously.
In recent years, differentiable rendering has emerged as a promising technique to reconstruct 3D contents for novel view synthesis [28, 3, 29, 56], surface reconstruction [48, 26, 23, 12], and animation [30, 35, 50]. In the field of edge reconstruction, several works introduced the idea of differentiable rendering to recover edge information [55, 20, 8]. Instead of detecting parametric 2D edge segments explicitly, they leverage pixel-wise 2D edge intensity images [45, 34] as a supervision signal and optimize a point-based 3D edge representation via differentiable multi-view rendering, resulting in an extracted 3D edge point set. Parametric 3D edges are then obtained by fitting lines and curves to the extracted 3D points. Since the edge-fitting step relies purely on the recovered 3D edge points but not the input 2D edge images, the fitted 3D edges may not align well with the 2D edge images. In addition, noise in the 3D edge points usually leads to discontinuous and fragmented edges.
To address these limitations, we propose SketchSplat, a differentiable edge reconstruction framework that directly optimizes parametric 3D edges by fitting them to the input 2D edge images using differentiable rendering. Our experimental results demonstrate that this leads to more accurate, complete, and compact reconstructions, more faithfully reflecting the original connectivity.
@@ 2 Related Work
The advancement of differentiable rendering opens up a new way for 3D reconstruction [28, 48, 23]. Inspired by NeRF [28], NEF [55] proposes to optimize an implicit neural radiance field with a multi-view rendering loss. After training, 3D edge points are extracted from the radiance field and curves are fitted to these points. NEAT [52] builds on VolSDF [53] and uses a signed distance field to represent both object surfaces and endpoints of 3D lines. The surface-based modeling approach restricts this method to objects with rich texture. EMAP [20] adopts an unsigned distance function (UDF), where sampled points can be displaced towards the zero-level set of the UDF to be more accurate. While methods based on neural field are time consuming to optimize, EdgeGS [8] represents 3D edge fields with explicit Gaussian points and leverages Gaussian splatting [18] to render and optimize Gaussian points, leading to more efficient edge reconstruction. A common issue of these methods is that the edge fitting process relies only on the extracted 3D edge points, hence noise in the reconstructed 3D points may lead to inaccurate and incomplete edges. Instead, we represent 3D edges as sketches defined parametrically by control points, scales, and opacity, and directly optimize these parameters via differentiable rendering. By leveraging sketch sampling and rasterization, our differentiable rendering aligns 3D edges with 2D edge images, resulting in more accurate and complete reconstructions.
@@
DiffVG [22] proposes a differentiable rasterizer that can optimize 2D vector graphics with 2D image supervision in a differentiable manner. However, the differentiation process relies on distance measurements between the rasterized curves and pixels, which needs to be custom tailored for each type of curve. 3Doodle [10] extends the idea of DiffVG [22] to reconstruct 3D curve-based object abstractions using multiple input views. They optimize a set of Bézier curves by minimizing the perceptual loss between the rasterized image and multi-view CLIP feature maps [38, 46]. Since CLIP features only contain abstract information, 3Doodle is not suitable for recovering accurate edges of objects like CAD models. Moreover, they approximate the projection of Bézier curves using an orthographic camera model. Instead, we get inspiration from 3DGS [18] and propose to optimize 3D sketches via sampling and rasterization of Gaussian points sampled from sketch parameters, which can be easily extended to different types of curves and achieves accurate 3D edge reconstruction.
@@ 4 Method
Given multi-view posed RGB images as input, our target is to recover the parametric edges for the 3D objects. Fig. 2 provides an overview of our approach. We firstly compute an edge image for each RGB image using our proposed 2D edge detector. Then we represent edges as 3D parametric sketches, including both lines $l\in\mathbb{R}^{2\times 3}$ and third-order Bézier curves with four 3D control points $c\in\mathbb{R}^{4\times 3}$. We initialize a set of sketches from existing edge reconstruction methods, and treat their control points, along with per-sketch scales and opacity, as optimizable parameters. During optimization, we use a sample-based method to convert sketches to Gaussian points, which can be rasterized onto 2D images to calculate the image loss. We back-propagate the gradient of the loss to optimize the sketch parameters. In addition, we introduce a set of topological operations to adaptively refine the sketches during training, yielding more accurate and compact edge reconstruction results.
Figure 2: Overview of SketchSplat pipeline: (a) We extract edge images using a novel approach (Sec. 4.1). (b) We then obtain initial edge points using EdgeGS [8] and (c) fit parametric sketches defined as lines and third-order Bézier curves. The core step of our approach is to optimize the sketches by (d) sampling and rasterizing them in a differentiable manner and back-propagating gradients of the image loss to (e) sketch parameters. In addition, during training we apply (f) a set of topological operations to improve compactness and connectivity of the reconstructed sketches. (g) After training, we take the optimized geometric parameters of sketches as the reconstructed 3D edges.
@@ 4.1 Edge Image Extraction
Previously, some methods [24, 52] extract edges with line detectors like LSD [47] and SOLD2 [31]. While these methods produce accurate parametric lines in 2D, they may fail to handle curve regions and result in incomplete results. Recent methods [55, 20, 8] use neural network-based edge detectors [45, 34] to detect edge intensity maps as supervision signal. However, since these detectors are not trained specifically on CAD-like objects, they may miss some significant edges in 2D or wrongly recover shading textures as real edges. Finally, we observed that these detectors introduce bias by consistently shifting detected edges slightly from certain real edges, which leads to multi-view inconsistency in 3D edge reconstruction (see supplementary). We address these issues by proposing a new edge detection method, which leverages geometric cues (e.g., per-pixel depth and normal estimation of 2D RGB images) to produce more accurate edge detection results.
For each RGB image, we combine information from three parts: foreground mask $A$, depth map $D$, and normal map $N$. The foreground mask can be extracted using the image alpha-channel (if available), a thresholding operation or possibly using semantic segmentation, providing object boundary information. We obtain a depth image for each view by reconstructing the scene with 2DGS [15] and rendering depth maps from each view. We use a normal estimator [54] to generate normal maps $N$ for each view. Then we calculate gradient maps $g(D)$ and $g(N)$ with the Sobel operator [43], threshold them with $t_{d}$ and $t_{n}$, and combine them via the OR operation, followed by a Gaussian filter $G_{f}(\cdot)$ to smooth the boundaries of the edges. In summary, we call our edge detector 2DGS-SN, defined as
@@ 4.2 Sketch Splatting
We represent edges as 3D parametric sketches, including both lines $l\in\mathbb{R}^{2\times 3}$ and third-order Bézier curves with control points $c\in\mathbb{R}^{4\times 3}$. This is the same as the output format of EdgeGS[8]. While EdgeGS optimizes 3D point positions first and then fits 3D lines and curves to the 3D points, we optimize the parameters of lines and curves directly via multi-view differentiable rasterization. In addition to the optimizable geometry parameters ($l$ or $c$), we equip each sketch with an optimizable opacity $o\in\mathbb{R}$ and a local scale $s\in\mathbb{R}^{3}$ to model the sketch intensity and thickness, respectively. These attributes will support Gaussian sampling and rasterization in the next steps.
@@
Our key idea is to bridge the gap between 2D edge images and 3D sketches via differentiable rendering and gradient descent optimization, which significantly improves sketch reconstruction over initialization method EdgeGS [8]. Inspired by 3DGS [18] and DiffVG [22], we choose 3D Gaussians as an intermediate representation to achieve differentiable sketch rendering.
For each gradient descent step, we first sample points from each sketch with a constant step size (5mm for all experiments), resulting in a point set $P$. Each point will be assigned the opacity and scale value of the corresponding sketch. To model the orientation, the main direction of each point is sampled as the local tangent direction of the sketch. Then we follow 3DGS[18] to rasterize the Gaussian points onto a chosen 2D view to generate a rendered edge image $E^{*}$. We calculate the $L_{1}$ loss between $E^{*}$ and the ground truth edge image $E_{gt}$, and back-propagate the gradient to update the sketch parameters $S_{i}=\{l_{i}\ \text{or}\ c_{i},o_{i},s_{i}\}$. We follow EdgeGS [8] to sample an equal number of foreground and background pixels for $L_{1}$ loss calculation.
@@ 4.3 Adaptive Topology Control
In our representation, each sketch has two end-points. If there exist two end-points, one from sketch $S_{i}$ and one from sketch $S_{j}$, that are sufficiently close (i.e. the point distance is smaller than threshold $th_{connect}=10mm$), we will connect sketch $S_{i}$ and $S_{j}$ by merging those two end-points into one point (see Fig. 2(f-i)). However, this means the number of optimizable points could differ from the number of control points $|\{l\}|*2+|\{c\}|*4$. To accommodate this, we maintain an optimizable point set $P$, and the control points of each line $l$ and curve $c$ are represented by indexing point set $P$. This representation also supports our additional topological operations.
@@
As shown in Fig. 2(f-ii), if one sketch is almost entirely covered by another, the smaller one will be merged into the larger one. Specifically, for sketch $S_{i}$ and $S_{j}$, we sample two point sets $P_{i}$ and $P_{j}$ and calculate the distance between each pair of points. Then for $P_{i}$, we calculate the ratio $r_{i\rightarrow j}$ of points that have close neighboring points in $P_{j}$, where the neighbor threshold is $th_{neighbor}=10mm$. If $r_{i\rightarrow j}>th_{overlap}=80\%$, sketch $S_{i}$ will be merged in sketch $S_{j}$. If both $r_{i\rightarrow j}$ and $r_{j\rightarrow i}$ surpass $th_{overlap}$, we will choose the sketch with larger overlapping ratio to be merged.
@@ 4.4 Implementation Details
Unless otherwise specified, we initialize with EdgeGS [8], which supplies a sufficient number of 3D edges to capture object structure. For each scene, we optimize the sketches for 1000 epochs with Adam. We update parameters only once for each epoch, where the loss is accumulated across all the training views. All of the scenes are scaled to fit within a $1\text{m}^{3}$ bounding box, and topology operation hyperparameters are set accordingly. Distance thresholds (e.g., $th_{connect}$) are set based on the desired precision and we set all of them as $1\%$ of the scene size ($1m^{3}$). The percentage and angular thresholds (e.g., $th_{overlap}$ and $th_{dir}$) work across different scenes without adjustment. All our experiments are implemented on a single RTX A5000 NVIDIA GPU. In average, each scene in ABC-NEF [55] dataset costs about 10 minutes for training, where both initialization and sketch optimization takes about 5 minutes, respectively.
@@ 5 Experiments
Table 1: Quantitative evaluation on ABC-NEF dataset [55]. Our method achieves the state-of-the-art performance on most metrics. A: Accuracy ($m$), C: Completeness ($m$), Rxx: Recall (threshold $x$ $m$), Pxx: Precision ($m$), Fxx: F-score ($m$).
Figure 3: Qualitative comparison on the ABC-NEF dataset [55]. Compared to the baselines, our method achieves superior accuracy and completeness, along with a favorable balance between completeness and compactness.
5.1 Experimental Setup
Datasets.
We evaluate our method on three datasets: ABC-NEF [55, 19] DTU [1], and Replica [44]. The ABC-NEF dataset [55, 19] comprises 115 CAD models, each containing 50 RGB images of the target object and a parametric edge file providing ground truth 3D edges. Following EMAP [20] and EdgeGS [8], we use 82 CAD models from ABC-NEF, excluding the ones with cylinders and spheres since smooth convex surfaces will produce silhouette edges that are not consistent in multi-views. For 2D edge image generation, we use PiDiNet [45], DexiNed [34], and our proposed 2DGS-SN as 2D edge detectors. The DTU dataset [1] contains multi-view images of real-world scenes. We follow EdgeGS [8] and evaluate on 6 DTU scenes. We use edge images and the pseudo ground truth edges provided by EMAP [20]. The Replica dataset [44] consists of multiple indoor scenes. We follow EdgeGS [8] to conduct qualitative evaluation on Replica [44].
Figure 4: Qualitative comparison on the DTU and Replica datasets [1, 44]. Enum is the edge number and edges are visualized with varied colors. Compared to the baselines, our method achieves higher compactness and retains faithful completeness.
Baselines.
We compare our method with five baseline methods, including the state-of-the-art 3D line mapping method (LIMAP [24]), and several methods based on volumetric differentiable rendering (NEF [55], NEAT [52], EMAP [20], EdgeGS [8]). Following EMAP and EdgeGS, we do not evaluate NEAT on the ABC-NEF dataset since NEAT often fails to train for texture-less objects.
Metrics.
Following the metrics used in baseline methods [20, 8], we evaluate Accuracy (A) and Completeness (C) in millimeters, and Recall (R), Precision (P), and F-score (F) as percentages at three thresholds ($5mm$, $10mm$, $20mm$). See supplementary material for more details.
5.2 Results
Figure 5: Ablation study on the key components of SketchSplat. Removing key components leads to noisier or redundant reconstructions.
Evaluation on ABC-NEF [55].
The qualitative and quantitative evaluation results on ABC-NEF dataset are shown in Fig. 3 and Tab. 1, respectively. Overall, the proposed method achieves a new state-of-the-art across most of the metrics. To the best of our knowledge, SketchSplat is the first method to achieve both accuracy (A) and completeness (C) below $7.0mm$, while also exceeding $90\%$ in recall (R5), precision (P5), and F-score (F5).
Tab. 1 shows LIMAP [24] achieves the highest accuracy due to its robust line mapping method. However, its completeness and recall are low, as LIMAP struggles with curve reconstruction and consistent edge detection results (Fig. 3(b)). NEF [55] produces more complete results in Fig. 3(c), but performs poorly on metrics as its fitting process is sensitive to noise in the extracted 3D points. EMAP [20] outperforms NEF with its unsigned distance field modeling and point shifting mechanism. However, Fig. 3(d) shows that EMAP still generates duplicated edges and misses some edges. With PidiNeT and DexiNed as 2D detector, EdgeGS [8] surpasses EMAP in recall and precision at $10mm$ and $20mm$ thresholds, but the trend reverses for measurements under $5mm$. This is due to the thickness of 2D edges introduces bias in smaller distance evaluations. When using our detector 2DGS-SN, EdgeGS performs better than EMAP to a large margin under $5mm$ metrics. As shown in Fig. 3, our SketchSplat achieves the highest precision and completeness compared to all the baselines.
As shown in Tab.1, the proposed 2DGS-SN significantly boosts the performance of EMAP, EdgeGS, and SketchSplat, addressing the bias from PiDiNet[45] and DexiNed [34]. Notably, across all edge detectors, our method consistently achieves higher Recall, Accuracy, and F-score than EdgeGS [8], demonstrating effectiveness of SketchSplat in aligning 3D edges with 2D edge images.
Method E-num Time (h)
NEF [55] 22.9 1:26
EMAP [20] 45.8 2:30
EdgeGS [8] 140.8 0:05
ours 44.3 0:10
Table 2: Comparison of the number of reconstructed edges (E-num) and training time on ABC-NEF dataset.
Evaluation on DTU [1] and Replica [44].
Fig. 4 presents qualitative comparisons on the DTU and Replica datasets. Compared to the baselines, our method consistently produces accurate and complete results using significantly fewer edges, demonstrating a strong balance between completeness and compactness. Although the ground-truth edges in the DTU scenes are not sufficiently accurate for reliable quantitative evaluation [8], our method still achieves metrics comparable to state-of-the-art methods EMAP and EdgeGS (see supplementary material for more details).
Speed Analysis.
Compactness.
EdgeGS achieves higher recall than NEF and EMAP due to its strict edge fitting process, which produces numerous short edges, covering more edge regions as the number of edges increases. Instead, our method incorporates adaptive topology control to reduce redundant edges, as shown in Tab. 2, Fig. 4, and Fig. 5(a,b,d). This mechanism helps SketchSplat achieve higher compactness, while maintaining the highest accuracy and completeness.
5.3 Ablation